Calculate the energy values of the electron in the second orbit of hydrogen
atom. What will be the radius of this orbit?
1
Expert's answer
2019-02-08T03:51:57-0500

1. To find the energy of an electron with principal quantum number n we should use the formula:

En=-(Z2* k)/n2

Where Z –atomic number

k- Bohr constant, k= 2.179*10-18 J

n- number of the energy level (principal quantum number)

For hydrogen atom Z=1, for an electron in the second orbit n=2

E2= -((1)2*2.179*10-18 J)/22 =-5.448*10-19 J

2. Bohr radius for nth orbit is calculated according to the formula:

rn = 0.529*n2/z

where n is a number of the orbit

z - atomic number

r2 = 0.529*22 /1 = 0.529*4 = 2.116 Angstrom

Answer: 1. r2 = 2.116 Angstrom, 2. E2= -5.448*10-19 J